# A New Formulation for Zeroth-Order Optimization of Adversarial EXEmples in Malware Detection

### v1

@@ 4.1 Experimental Results
We present the comparison between ZEXE, GAMMA, OZD, and RS against the baseline models MalConv (Table 2) and GBDT-EMBER (Table 3).
Table 2: Evasion rate, average and standard deviation of the number of queries required to reach a confidence $<0.5$ on MalConv computed on $200$ malware.
Algorithm ER $\mu_{\text{queries}}$ $\sigma_{\text{queries}}$ $\mu_{\text{size}}$ $\sigma_{\text{size}}$
@@
GAMMA $[\lambda=10^{-3}]$ $0.14$ $1.64$ $1.26$ $632.278$ $829.358$
GAMMA $[\lambda=10^{-5}]$ $0.24$ $12.80$ $29.80$ $648.923$ $836.814$
GAMMA $[\lambda=10^{-7}]$ $0.58$ $14.38$ $29.24$ $881.529$ $860.288$
@@
In Table 2, we observe that ZEXE achieves the highest evasion rate against MalConv. This indicates that the adversarial EXEmples crafted by our algorithm are more effective than those produced by the state-of-the-art approach, GAMMA. Also, ZEXE is able to achieve this results with minimal amount of injected bytes: due to different structures of executable and different constraints, the mean perturbation achieved is 211KB, against the 881KB achieved by GAMMA. Hence, ZEXE only needs only one fourth of the perturbation needed by GAMMA. However, we remark that ZEXE requires a higher number of queries than the other algorithms to reach evasion. This is mainly due to the fact that exploration is performed through a random search strategy, while the higher evasion rate is justified by the local exploration methods used. Note also that for MalConv a simple finite-difference strategy as OZD with our section injection obtains an high evasion rate with few queries required.
Table 3: Evasion rate, average and standard deviation of the number of queries required to reach a confidence $<0.8$ on GBDT-EMBER computed on $200$ malware.
Algorithm ER $\mu_{\text{queries}}$ $\sigma_{\text{queries}}$ $\mu_{\text{size}}$ $\sigma_{\text{size}}$
@@
GAMMA $[\lambda=10^{-3}]$ $0.08$ $2.56$ $4.12$ $1261.796$ $216.930$
GAMMA $[\lambda=10^{-5}]$ $0.12$ $3.08$ $6.26$ $1258.171$ $211.115$
GAMMA $[\lambda=10^{-7}]$ $0.20$ $10.05$ $18.52$ $1513.767$ $404.718$
@@
In Table 3, we observe that ZEXE once again achieves the highest evasion rate with a small average perturbation size against GBDT-EMBER. In this case, the difference between attacks is deepened by the difference in evasion rate (ER). In particular, ZEXE computes 17% successfully-evasive adversarial EXEmples more than the best competing version of GAMMA. Regarding the manipulation size, ZEXE only needs one third of the manipulation injected by GAMMA, remarking the superiority of our novel attack. Additionally, we note that RS matches the performance of GAMMA, implying that (i) the choice of the initial benignware to is crucial for the success of the attack; and (ii) that our choice of leveraging only printable strings inside the manipulation had boosted the effectiveness of ZEXE. Lastly, OZD performs worse against GBDT-EMBER than the other approaches. This is likely due to the lack of an exploration strategy and the non-convexity of the target function, which prevents the algorithm from finding good solutions. Hence, we can confirm that ZEXE mitigates this problem through the exploration procedures we included.
Table 4: Average $\mu_{\text{conf}}$ and standard deviation $\sigma_{\text{conf}}$ of the confidence $F$ of MalConv and GBDT-EMBER computed on on $200$ adversarial EXEmples.
@@
GAMMA $[\lambda=10^{-3}]$ $0.8540$ $0.2966$ $0.9344$ $0.1960$
GAMMA $[\lambda=10^{-5}]$ $0.7614$ $0.3635$ $0.9001$ $0.2538$
GAMMA $[\lambda=10^{-7}]$ $0.4382$ $0.4427$ $0.8452$ $0.3010$
@@
Finally, we report in Table 4 the mean and standard deviation of the confidence both MalConv and GBDT-EMBER attributed to the adversarial EXEmples crafted by the selected attacks. ZEXE diminishes more the average confidence of both classifiers on all the adversarial EXEmples it crafts, w.r.t. all its competitors. While GAMMA and OZD manages to rival our attack on MalConv, where the difference is roughly 5% , on GBDT-EMBER we assist to a reduction of more than 11% of the average confidence.
@@ 5 Related Work
Formulation of Adversarial EXEmples. As already described in Sec. 2, [12] propose a formulation for computing adversarial EXEmples through the introduction of practical manipulations inside the optimization problem. The authors also show that most of the proposed attacks can be recasted inside their proposal, highlighting its generality. [27] propose a general formulation for adversarial evasion attacks that can be customized to deal with Windows programs as well. To do so, the authors include an “oracle” function that informs whether manipulations maintain the original functionality or not. However, unlike our work, both proposals can not be used to formulate theoretical guarantees of optimization algorithms used to minimize the optimization problem. Also, while we do not need any “oracle” since we deal with manipulations which preserve functionality by design, [27] introduce such a function but they do not explain how to practically apply it inside the optimization problem, rendering it unapplicable.
Gradient-free Adversarial EXEmples attacks. The state of the art provides two different approaches to create adversarial EXEmples: (i) algorithms that optimize the sequence of manipulations [2, 34]; and (i) genetic algorithms [12, 13] that fine-tune injected content inside samples. Among them, [12] propose GAMMA, that leverages section injection to include benignware-extracted content inside malware, controlling the perturbation size with a regularization parameter.
@@
However, all these techniques lack theoretical guarantees, binding the results of attacks to mostly-random heuristics with suboptimal performances. Also, these attacks either require resource-demanding sandbox validation [8, 34], or they can not fix the perturbation size a priori due to the nature of the chosen practical manipulation [2, 12]. Lastly, GAMMA, which is one of the mostly-used attack, strictly depends on a distribution of benignware chosen a priori, thus heavily influencing the evasion rate. On the contrary, our formulation of zero-order attacks cover all these issues. On top of this, we prove theoretical guarantees of ZEXE, and we show its superiority in terms of trade-off between queries, perturbation. Lastly, thanks to its adaptive formulation, ZEXE requires minimal hyper-parameter tuning, and it can be applied against any Windows malware detector.
@@
Zeroth-order optimization algorithms. In the literature many methods were proposed to tackle optimization problem using only function values and finite-difference methods have been widely studied and analyzed in different settings with different gradient approximations [26, 15, 33, 3, 16, 30, 31, 22, 6, 7]. Some of these [7, 6] have also been applied to compute adversarial examples of image and audio samples as well. [7] introduce a finite-difference method that build linear interpolation gradient estimator [3] and exploits sparsity by including a constraints on the approximation. However, due to the domain constraints imposed by the Windows PE file format, this method cannot be used for solving this problem directly, and, due to the fixed perturbation size applied by practical manipulation, it is not evident that this algorithm can exploit sparsity in this domain.
On the contrary, our formalization allows to use these methods as we did for OZD in Sec. 4. Moreover, both empirical and theoretical evidence suggest that ZEXE, our new optimization algorithm, perform better than unstructured methods [3, 9]. Also, ZEXE avoids stagnation in stationary points thanks to the exploration we included, differently from other finite-difference methods, and we empirically see the effect of this improvement in Sec. 4.1, achieving higher evasion rate than structured zeroth-order method OZD.
@@ 6 Contributions, Limitations, and Future Work
In this work, we have proposed a novel formulation for computing adversarial EXEmples, carefully-manipulated Windows programs that bypass detection, that can enable the application of sound and efficient gradient-free optimization algorithms even in the presence of complex domain constraints. As a by-product, we define ZEXE, a zeroth-order technique that not only is backed-up by theoretical guarantees, but also drastically improves adversarial EXEmples in terms of both evasion rate and injected content.
@@
However, we only investigated the effect of one practical manipulation $\rho$, while more have been proposed [12]. Nevertheless, our formulation is general enough to tackle all of them, by repeating the same experiments with different implementations of $\rho$. We tested ZEXE against GAMMA and OZD, removing other strategies from the comparison. While this can be accomplished through more experiments, we want to remark that [34, 8, 2] need either samples to be validated inside isolated environment, or training complex reinforcement learning policies thus increasing the complexity of their setup which is absent for ZEXE. Also, ZEXE needs, on average, more queries than its competitors to compute successfully-evasive adversarial EXEmples. However, as we already discussed in Sec. 4.1, ZEXE achieves the best trade-off between the superior evasion rate and the perturbation size, with GAMMA and OZD failing in both.
@@
We want to remark that our work, even if focused on evasion attacks, has the purpose of expanding the robustness testing techniques in this domain, which present inherent complexities that must be tackled accordingly. Thus, we expect our work to be used not as an offensive strategy, but as a testing technique applied before placing detectors in productions to avoid the spreading of weak detectors against adversarial EXEmples.
@@
We conclude by envisioning that our new formulation will potentially enable more investigation of sound and theoretically-grounded optimization algorithms in the domain of adversarial EXEmples. Due to the easiness of application and hyper-parameter tuning, we envision ZEXE to be used to benchmark the robustness of Windows malware detectors with different manipulation functions. This would produce a robustness leader board, similarly to what has been done in the image classification domain [10], which can be used to fairly compare the performance of all the proposed detectors in the literature, in particular the ones developed to resist adversarial EXEmples attacks [25, 19, 18].
@@
Acknowledgments. L. R. acknowledges the financial support of the European Research Council (grant SLING 819789), the European Commission (ELIAS 101120237), the US Air Force Office of Scientific Research (FA8655-22-1-7034), the Ministry of Education, University and Research (grant ML4IP R205T7J2KP; grant BAC FAIR PE00000013 funded by the EU - NGEU) and the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF-1231216. M. R. acknowledges the financial support of the European Research Council (grant SLING 819789). M. R. is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). F.R. and L.D. acknowledge the financial support of project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU; project FAIR (PE00000013) under the NRRP MUR program funded by the EU - NGEU. This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains.
@@ Appendix A Auxiliary Results
In this appendix, we state every lemma and proposition required to prove the theoretical results of Sec. 3.3.
Notation. In the following appendices, we denote with $\mathcal{F}_{k}$ the filtration $\sigma(G_{0},\cdots,G_{k-1})$. We denote with $F_{z}$ the target function of Eq. 5 for an arbitrary malware $z\in Z$ and $F_{h}$ as its smooth approximation defined as
@@
where $\mathbb{B}^{d}:=\{x\in\mathbb{R}^{d}\,|\,\|x\|\leq 1\}$ is the $d$ dimensional unit-ball. Note that, to improve the readability of lemmas and proofs, we omit the indication of $z$ in $F_{h}$.
Lemma 1 (Square norm of sum of orthogonal vectors).
Let $(v_{i})_{i=1}^{\ell}\in\mathbb{R}^{d}$ such that for $i,j=1,\cdots,\ell$ and $i\neq j$,
$v_{i}^{\intercal}v_{i}=1\qquad\text{and}\qquad v_{i}^{\intercal}v_{j}=0.$
Then
$\Bigg{\|}\sum\limits_{i=1}^{\ell}v_{i}\Bigg{\|}^{2}=\sum\limits_{i=1}^{\ell}\|v_{i}\|^{2}$
@@
By definition,
$\Bigg{\|}\sum\limits_{i=1}^{\ell}v_{i}\Bigg{\|}^{2}=\langle\sum\limits_{i=1}^{\ell}v_{i},\sum\limits_{i=1}^{\ell}v_{i}\rangle.$
By linearity,
$\Bigg{\|}\sum\limits_{i=1}^{\ell}v_{i}\Bigg{\|}^{2}=\sum\limits_{i=1}^{\ell}\sum\limits_{j=1}^{\ell}\langle v_{i},v_{j}\rangle.$
Splitting the sum,
$\Bigg{\|}\sum\limits_{i=1}^{\ell}v_{i}\Bigg{\|}^{2}=\sum\limits_{i=1}^{\ell}\sum\limits_{j=i}\langle v_{i},v_{j}\rangle+\sum\limits_{i=1}^{\ell}\sum\limits_{j\neq i}\langle v_{i},v_{j}\rangle=\sum\limits_{i=1}^{\ell}\langle v_{i},v_{i}\rangle+\sum\limits_{i=1}^{\ell}\sum\limits_{j\neq i}\langle v_{i},v_{j}\rangle.$
By orthogonality, we have $\langle v_{i},v_{j}\rangle=0$ for $i\neq j$. Thus,
$\Bigg{\|}\sum\limits_{i=1}^{\ell}v_{i}\Bigg{\|}^{2}=\sum\limits_{i=1}^{\ell}\langle v_{i},v_{i}\rangle=\sum\limits_{i=1}^{\ell}\|v_{i}\|^{2}.$
Proposition 1 (Smoothing properties).
Let $F_{h}$ be the smooth approximation defined in Eq. 11. Then, if $F_{z}$ is $L$-Lipschitz continuous - i.e. $\forall x,y\in\mathbb{R}^{d}$, $|F_{z}(x)-F_{z}(y)|\leq L\|x-y\|$, $F_{h}$ is $L$-Lipschitz continuous, differentiable and for every $x,y\in\mathbb{R}^{d}$
$\|\nabla F_{h}(x)-\nabla F_{h}(y)\|\leq\frac{L\sqrt{d}}{h}\|x-y\|\quad\text{and}\quad F_{h}(x)\leq F_{z}(x)+Lh.$
Proof.
This is a standard result proposed and proved in different works - see for example [14, Lemma 8],[23, Proposition 2.2] and [36]. ∎
@@
The proof follows the same line of [31, Lemma 1]. ∎
@@
$\mathbb{E}_{G_{k}}[F_{h_{k}}(v_{k+1})\,|\,\mathcal{F}_{k}]\leq F_{h_{k}}(v_{k})-\gamma_{k}\|\nabla F_{h_{k}}(v_{k})\|^{2}+\frac{L\sqrt{d}}{h_{k}}\gamma_{k}^{2}\mathbb{E}_{G_{k}}[\|g_{k}(v_{k})\|^{2}\,|\,\mathcal{F}_{k}].$
@@
In the following, we denote with $g_{k}(\cdot):=g_{(G_{k},h_{k})}(\cdot)$. By Proposition 1, we have that $F_{h_{k}}$ is $L\sqrt{d}/h_{k}$ smooth. Thus, by the Descent Lemma [28],
$F_{h_{k}}(v_{k+1})\leq F_{h_{k}}(v_{k})-\gamma_{k}\langle\nabla F_{h_{k}}(v_{k}),g_{k}(v_{k})\rangle+\frac{L\sqrt{d}}{h_{k}}\gamma_{k}^{2}\|g_{k}(v_{k})\|^{2}.$
@@
$\mathbb{E}_{G_{k}}[F_{h_{k}}(v_{k+1})\,|\,\mathcal{F}_{k}]\leq F_{h_{k}}(v_{k})-\gamma_{k}\|\nabla F_{h_{k}}(v_{k})\|^{2}+\frac{L\sqrt{d}}{h_{k}}\gamma_{k}^{2}\mathbb{E}_{G_{k}}[\|g_{k}(v_{k})\|^{2}\,|\,\mathcal{F}_{k}].$
Theorem A.1 (Non-convex Non-smooth Rates).
Under Assumption 1, let $(v_{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then,
$\sum\limits_{i=0}^{k}\gamma_{i}\mathbb{E}[\|\nabla F_{h_{i}}(v_{i})\|^{2}]\leq F_{h_{0}}(v_{0})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=0}^{k}\frac{\gamma_{i}^{2}}{h_{i}}.$
Proof.
By Lemma 4 and Lemma 2, we get
$\mathbb{E}_{G_{k}}[F_{h_{k}}(v_{k+1})\,|\,\mathcal{F}_{k}]\leq F_{h_{k}}(v_{k})-\gamma_{k}\|\nabla F_{h_{k}}(v_{k})\|^{2}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\frac{\gamma_{k}^{2}}{h_{k}}.$
Taking the full expectation,
$\gamma_{k}\mathbb{E}[\|\nabla F_{h_{k}}(v_{k})\|^{2}]\leq\mathbb{E}[F_{h_{k}}(v_{k})-F_{h_{k}}(v_{k+1})]+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\frac{\gamma_{k}^{2}}{h_{k}}.$
Summing from $i=0,\cdots,k$,
$\sum\limits_{i=0}^{k}\gamma_{i}\mathbb{E}[\|\nabla F_{h_{i}}(v_{i})\|^{2}]\leq\mathbb{E}[F_{h_{0}}(v_{0})-F_{h_{k}}(v_{k+1})]+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=0}^{k}\frac{\gamma_{i}^{2}}{h_{i}}.$
By the definition of $F_{h_{k}}$, we have $F_{h_{k}}(v)\geq\min F_{z}$ for every $v\in\mathbb{R}^{d}$ and $h_{k}>0$. Thus,
$\sum\limits_{i=0}^{k}\gamma_{i}\mathbb{E}[\|\nabla F_{h_{i}}(v_{i})\|^{2}]\leq F_{h_{0}}(v_{0})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=0}^{k}\frac{\gamma_{i}^{2}}{h_{i}}.$
This concludes the proof.

### v2

@@ 4.3 Experimental Results
We first present the comparison between ZEXE, GAMMA, OZD, and RS against the baseline models MalConv (Table 3) and GBDT-EMBER (Table 4).
Table 3: Evasion rate achieved against MalConv with how many queries queries ($\mu_{\text{queries}}$, $\sigma_{\text{queries}}$) and size ($\mu_{\text{KB}}$ / $\sigma_{\text{KB}}$) computed on $200$ samples.
Algorithm ER $\mu_{\text{queries}}$ $\sigma_{\text{queries}}$ $\mu_{\text{size}}$ (KB) $\sigma_{\text{size}}$ (KB)
@@
GAMMA $\lambda=10^{-3}]$ $0.14$ $1.64$ $1.26$ $632.278$ $829.358$
GAMMA $\lambda=10^{-5}]$ $0.24$ $12.80$ $29.80$ $648.923$ $836.814$
GAMMA $\lambda=10^{-7}]$ $0.58$ $14.38$ $29.24$ $881.529$ $860.288$
@@
In Table 3, we observe that ZEXE achieves the highest evasion rate against MalConv. This indicates that the adversarial EXEmples crafted by our algorithm are more effective than those produced by the state-of-the-art approach, GAMMA. Also, ZEXE is able to achieve these results with a minimal amount of injected bytes: due to different structures of executable and different constraints, the mean perturbation achieved is 211KB, against the 881KB achieved by GAMMA. Hence, ZEXE only needs one fourth of the perturbation needed by GAMMA. However, we remark that ZEXE requires a higher number of queries than the other algorithms to reach evasion. This is mainly due to the fact that exploration is performed through a random search strategy, while the higher evasion rate is justified by the local exploration methods used. Note also that for MalConv, a simple finite-difference strategy as OZD with our section injection obtains a high evasion rate with few queries required.
Table 4: Evasion rate achieved against GBDT-EMBER with how many queries queries ($\mu_{\text{queries}}$, $\sigma_{\text{queries}}$) and size ($\mu_{\text{KB}}$ / $\sigma_{\text{KB}}$) computed on $200$ samples.
Algorithm ER $\mu_{\text{queries}}$ $\sigma_{\text{queries}}$ $\mu_{\text{size}}$ (KB) $\sigma_{\text{size}}$ (KB)
@@
GAMMA $\lambda=10^{-3}]$ $0.08$ $2.56$ $4.12$ $1261.796$ $216.930$
GAMMA $\lambda=10^{-5}]$ $0.12$ $3.08$ $6.26$ $1258.171$ $211.115$
GAMMA $\lambda=10^{-7}]$ $0.20$ $10.05$ $18.52$ $1513.767$ $404.718$
@@
In Table 4, we observe that ZEXE once again achieves the highest evasion rate with a small average perturbation size against GBDT-EMBER. In this case, the difference between attacks is deepened by the difference in evasion rate (ER). In particular, ZEXE computes 17% successfully-evasive adversarial EXEmples more than the best competing version of GAMMA. Regarding the manipulation size, ZEXE only needs one third of the manipulation injected by GAMMA. Additionally, we note that RS matches the performance of GAMMA, implying that (i) the choice of the initial benignware is crucial for the success of the attack; and (ii) that our choice of leveraging only printable strings inside the manipulation had boosted the effectiveness of ZEXE. Lastly, OZD performs worse against GBDT-EMBER than the other approaches. This is likely due to the lack of an exploration strategy and the non-convexity of the target function, which prevents the algorithm from finding good solutions.
Table 5: Average $\mu_{\text{conf}}$ and standard deviation $\sigma_{\text{conf}}$ of the confidence $F$ of MalConv and GBDT-EMBER computed on $200$ adversarial EXEmples.
@@
GAMMA $\lambda=10^{-3}]$ $0.8540$ $0.2966$ $0.9344$ $0.1960$
GAMMA $\lambda=10^{-5}]$ $0.7614$ $0.3635$ $0.9001$ $0.2538$
GAMMA $\lambda=10^{-7}]$ $0.4382$ $0.4427$ $0.8452$ $0.3010$
@@
In Table 5, we report the mean and standard deviation of the confidence both MalConv and GBDT-EMBER attributed to the adversarial EXEmples crafted by the selected attacks. ZEXE diminishes more the average confidence of both classifiers on all the adversarial EXEmples it crafts, w.r.t. all its competitors. While GAMMA and OZD manage to rival our attack on MalConv, where the difference is roughly 5%, on GBDT-EMBER we observe a reduction of more than 11% of the average confidence.
Finally, we consider the second setting where high-dimensional perturbations are used. In Table 6, we report the evasion rate, the average, and standard deviation of the perturbation size of GAMMA and variants of ZEXE. As we can observe, due to the high-dimensionality, exploration strategies like random search and Langevin achieve worse performance than GAMMA, suggesting that a more sophisticated method should be used. Instead, performing the exploration steps with genetic algorithm resulted to be effective, and it allowed for a higher evasion rate than GAMMA even when a higher number of sections is used.
Table 6: Evasion rate of GBDT-EMBER achieved on $700$ malware.
Algorithm ER $\mu_{\text{size}}$ (MB) $\sigma_{\text{size}}$ (MB)
GAMMA ($30$ sections) $0.4743$ $22.394$ $8.389$
GAMMA ($40$ sections) $0.5354$ $32.141$ $11.172$
GAMMA ($50$ sections) $0.5100$ $36.214$ $11.562$
ZEXE-RS $0.2968$ $11.113$ $3.017$
ZEXE-LAN $0.3271$ $11.113$ $3.017$
ZEXE-GA 0.5974 11.113 3.017
@@ 5 Related Work
Formulation of Adversarial EXEmples. As already described in Sec. 2, [12] proposes a formulation for computing adversarial EXEmples through the introduction of practical manipulations inside the optimization problem. The authors also show that most of the proposed attacks can be recast inside their proposal, highlighting its generality. [29] propose a general formulation for adversarial evasion attacks that can be customized to deal with Windows programs as well. To do so, the authors include an “oracle” function that informs whether manipulations maintain the original functionality or not. Also, while we do not need any “oracle” since we deal with manipulations which preserve functionality by design, [29] introduces such a function, but they do not explain how to practically apply it inside the optimization problem, rendering it unapplicable.
@@
Gradient-free Adversarial EXEmples attacks. The state of the art provides two different approaches to create adversarial EXEmples: (i) algorithms that optimize the sequence of manipulations [1, 37]; and (i) genetic algorithms [12, 14] that fine-tune injected content inside samples. Among them, [12] proposes GAMMA, which leverages section injection to include benignware-extracted content inside malware, controlling the perturbation size with a regularization parameter. However, all these techniques lack theoretical guarantees, binding the results of attacks to mostly-random heuristics with suboptimal performances even for synthetic optimization problems or in controlled scenarios, without any theoretical support to describe their behavior. Also, these attacks either require resource-demanding sandbox validation [9, 37], or they can not fix the perturbation size a priori due to the nature of the chosen practical manipulation [1, 12]. Lastly, GAMMA, one of the most-used attacks, strictly depends on a distribution of a pre-selected distribution of benignware, thus heavily influencing the evasion rate. On the contrary, our formulation of zeroth-order attacks covers all these issues. On top of this, we prove theoretical guarantees of ZEXE to support and describe the algorithm behavior under classical optimization hypotheses, and we show that it provides better performance in terms of trade-off between queries and perturbation with respect to other competing approaches.
@@
Zeroth-order Optimization Algorithms. In the literature, many methods were proposed to tackle the optimization problem using only function values, and finite-difference methods have been widely studied and analyzed in different settings with different gradient approximations [28, 16, 36, 3, 17, 33, 32, 23, 7, 8]. Some of these [8, 7] have also been applied to compute adversarial examples of image and audio samples as well. [8] introduces a finite-difference method that builds a linear interpolation gradient estimator [3] and exploits sparsity by including constraints on the approximation. However, due to the domain constraints imposed by the Windows PE file format, this method cannot be used for solving this problem directly, and, due to the fixed perturbation size applied by practical manipulation, it is not evident that this algorithm can exploit sparsity in this domain. On the contrary, our formalization allows us to use these methods as we did for OZD in Sec. 4. Moreover, both empirical and theoretical evidence suggest that ZEXE, our new optimization algorithm, performs better than unstructured methods [3, 10] in controlled scenarios. Also, ZEXE avoids stagnation in stationary points thanks to the exploration we included, differently from other finite-difference methods, and we empirically see the effect of this improvement in Sec. 4.3, achieving a higher evasion rate than the structured zeroth-order method OZD.
@@ 6 Contributions, Limitations, and Future Work
In this work, we have proposed a novel formulation for computing adversarial EXEmples that can enable the application of sound and efficient gradient-free optimization algorithms even in the presence of complex domain constraints. Thus, we define ZEXE, a zeroth-order technique that allows for better performance in terms of the evasion rate and size of the injected content.
@@
However, we only investigated the effect of one practical manipulation $\rho$, while more have been proposed [12]. Nevertheless, our formulation is general enough to tackle all of them by repeating the same experiments with different implementations of $\rho$. We tested ZEXE against GAMMA and OZD, removing other strategies from the comparison. While this can be accomplished through more experiments, we want to remark that [37, 9, 1] need either samples to be validated inside an isolated environment, or training complex reinforcement learning policies, thus increasing the complexity of their setup, which is absent for ZEXE.
Also, ZEXE needs, on average, more queries than its competitors to compute successfully-evasive adversarial EXEmples, thus being potentially detectable. Such is caused by both the exploration strategy (random search requires many iterations before sampling a good perturbation due to the dimensionality of the perturbation space) and the finite-difference approximation of the gradient. However, this requirement could potentially be reduced in future work by adopting more sophisticated exploration strategies and by also approximating second-order information.
Moreover, we did not assess the impact of the initialization of adversarial perturbation in terms of convergence and stability of ZEXE (i.e., initialize all values with zero, or content harvested from benign software [12]). Although we limited our analysis to the usage of printable characters, we observed in Sec. 4.3 that ZEXE still achieves the best trade-off between evasion rate and perturbation size with respect to the considered approaches.
@@
We want to remark that our work, even if focused on evasion attacks, has the purpose of expanding the robustness testing techniques in this domain, which present inherent complexities that must be tackled accordingly. Thus, we expect our work to be used not as an offensive strategy, but as a testing technique applied before placing detectors in production to avoid the spreading of weak detectors against adversarial EXEmples.
@@
We conclude by envisioning that our new formulation will potentially enable more investigation of sound and theoretically-grounded optimization algorithms in the domain of adversarial EXEmples. Due to the ease of application and hyper-parameter tuning, we envision ZEXE to be used to benchmark the robustness of Windows malware detectors with different manipulation functions. This would produce a robustness leader board, similarly to what has been done in the image classification domain [11], which can be used to fairly compare the performance of all the proposed detectors in the literature, in particular the ones developed to resist adversarial EXEmples attacks [26, 20, 19]. Such would require a massive experimental analysis comprising more detectors and larger test datasets, providing a complete and comprehensive comparison between approaches on a large-scale setup. Also, ZEXE could be used to perform security evaluations of commercial products, and we envision future analysis to consider real-world antivirus programs as well [12].
Acknowledgments
@@
This work was partially supported by projects SERICS (PE00000014) and FAIR (PE00000013) under the NRRP MUR program funded by the EU - NGEU; and by European Research Council (grant SLING 819789); and by the European Commission (Horizon Europe grant ELIAS 101120237); and by the US Air Force Office of Scientific Research (FA8655-22-1-7034); and by the Ministry of Education, University and Research (FARE grant ML4IP R205T7J2KP); and by a France 2030 support managed by the Agence Nationale de la Recherche, under the reference ANR-23-PEIA-0004 (PDE-AI project); and by “InfoAICert” funded by the MUR program “Fondo italiano per le scienze applicate” (FISA-2023-00128). This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains.
@@ Appendix A Auxiliary Results
We report here every lemmas and propositions required to prove the theoretical results of Sec. 3.4.
Notation. From now on, we denote with $\mathcal{F}_{k}$ the filtration $\sigma(G_{0},\cdots,G_{k-1})$. We denote with $F_{z}$ the target function of Eq. 5 for a malware $z\in Z$ and $F_{h}$ as its smooth approximation defined as
@@
where $\mathbb{B}^{d}:=\{x\in\mathbb{R}^{d}\,|\,\|x\|\leq 1\}$ is the $d$ dimensional unit-ball. To improve the readability of lemmas and proofs, we omit the indication of $z$ in $F_{h}$.
Proposition 1 (Smoothing properties).
Let $F_{h}$ be the smooth approximation defined in Eq. 11. Then, if $F_{z}$ is $L$-Lipschitz continuous - i.e., $\forall x,y\in\mathbb{R}^{d}$, $|F_{z}(x)-F_{z}(y)|\leq L\|x-y\|$, $F_{h}$ is $L$-Lipschitz continuous, differentiable and for every $x,y\in\mathbb{R}^{d}$
$\|\nabla F_{h}(x)-\nabla F_{h}(y)\|\leq\frac{L\sqrt{d}}{h}\|x-y\|,$
and
$F_{h}(x)\leq F_{z}(x)+Lh.$
@@
This is a standard result proposed and proved in different works - see for example [15, Lemma 8],[25, Proposition 2.2] and [39]. ∎
@@
The proof follows the same line of [32, Lemma 1]. ∎
@@
$\mathbb{E}_{G_{k}}[F_{h}(v_{k+1})\,|\,\mathcal{F}_{k}]$
$\leq F_{h}(v_{k})-\gamma_{k}\|\nabla F_{h}(v_{k})\|^{2}+\frac{L\sqrt{d}}{h}\gamma_{k}^{2}\mathbb{E}_{G_{k}}[\|g_{k}(v_{k})\|^{2}\,|\,\mathcal{F}_{k}].$
@@
In the following, we denote with $g_{k}(\cdot):=g_{(G_{k},h)}(\cdot)$. By Proposition 1, we have that $F_{h}$ is $L\sqrt{d}/h$ smooth. Thus, by the Descent Lemma [30],
$F_{h}(v_{k+1})$
$\leq F_{h}(v_{k})-\gamma_{k}\langle\nabla F_{h}(v_{k}),g_{k}(v_{k})\rangle+\frac{L\sqrt{d}}{h}\gamma_{k}^{2}\|g_{k}(v_{k})\|^{2}.$
@@
$\mathbb{E}_{G_{k}}[F_{h}(v_{k+1})\,|\,\mathcal{F}_{k}]$
$\leq F_{h}(v_{k})-\gamma_{k}\|\nabla F_{h}(v_{k})\|^{2}+\frac{L\sqrt{d}}{h}\gamma_{k}^{2}\mathbb{E}_{G_{k}}[\|g_{k}(v_{k})\|^{2}\,|\,\mathcal{F}_{k}].$